# Effective and Evasive Fuzz Testing-Driven Jailbreaking Attacks against LLMs

### v1

@@ 29                   $S\leftarrow S\cup\{T\}$
Algorithm 1 Jailbreak Prompts Generation Procedure
Specifically, our goals are: 1) To circumvent the safety mechanisms of the target LLM, ensuring that it responds to a harmful query instead of opting not to answer. 2) To ensure that the responses to these questions are not only correct and relevant but also contain the harmful content intended by the query.
4 Methodology of our method
In this section, we first present the general attack framework of our method and then describe its key components.
4.1 Framework of our method
our method is an automated black-box jailbreaking attack method based on fuzz testing, which encompasses a cyclical regimen of seed initialization, seed selection, mutation, and execution.
  • Seed initialization. The first step is to generate or select initial inputs (seeds) from the current pool. We confront a bifurcated decision: leveraging pre-existing, human-crafted jailbreaking templates [55] or initiating seeds stochastically [27]. Human-written templates, however, often lack scalability and adaptability, making it difficult and time-consuming to find a suitable one, especially for new questions with limited references. Conversely, randomly initialized jailbreaking templates can result in meaningless sequences, making it challenging to create semantically coherent jailbreaks. To overcome these limitations, our method initializes with an empty seed pool, eliminating the need for existing jailbreaking templates. To address the initial challenge of limited seed availability, we introduce a pre-jailbreak phase where each problem undergoes several initial jailbreak attempts (e.g., 5 times). Successful jailbreaks from this phase are bypassed to conserve the query budget. Thus, the overall process in our method consists of two stages: the pre-jailbreak and final-jailbreak stages. This dual-phase strategy enhances the efficiency and effectiveness of our method.
  • Seed selection. Seed selection chooses specific seeds that are likely to achieve the attack goals to undergo mutation. There exist various seed selection strategies, such as random selection, Round Robin, UCB, and MCTS-Explore [55]. Follow [55], we use MCTS-Explore in our method by default.
  • Mutation. These selected seeds undergo systematic modification through mutation to create new jailbreaking templates. Traditional mutation strategies employed by fuzzers [^10] https://lcamtuf.coredump.cx/afl/ are primarily designed for binary or structured data. Directly applying these strategies to natural language inputs can lead to syntactically incorrect or semantically nonsensical inputs. In addition, various studies [31, 55] utilize specific mutation operators, such as “generate", “crossover", “expand", “shorten", and “rephrase", to generate mutations based on existing human-written jailbreaking templates. However, since these mutation methods are closely tied to the initial templates and are not specifically adapted to the harmful question, they often produce many redundant and ineffective mutations. In our method, we design three novel question-dependent mutation strategies capable of generating mutations with semantic meaning and limited token length by leveraging the power of an LLM helper. As a result, we can significantly improve attack performance and reduce attack costs.
  • Execution. Finally, the mutated seeds are evaluated using a judge model to distinguish between successful and unsuccessful jailbreaks. Successful jailbreaking templates are retained in the seed pool, while unsuccessful ones are discarded. To determine whether the response indicates a successful jailbreak, attackers can employ either human annotators [33, 57] or LLMs [32, 44, 47] for labeling assistance. However, using human annotators is not scalable and impractical for automatic fuzzing. Additionally, LLMs also suffer from inaccuracies. In our method, we implement a two-level automatic judge module to distinguish genuine successful jailbreaks, further decreasing query costs to victim LLMs.
Please refer to the Appendix for more detailed explanations of certain terminologies. The overview of our method is shown in Figure 2, and the overall algorithm is detailed in Algorithm 1.
@@ 4.2 Template Mutation and Optimization
Mutation schemes. Unlike existing methods that demand extensive manual effort and struggle with generalization, our method uses a LLM (e.g., GPT-3.5 turbo), as a mutator to automatically create jailbreaking templates for specific questions.
In the pre-jailbreak stage, we design Role-play and Contextualization mutation operators to create customized jailbreaking templates tailored to each question. Successful templates from this phase are incorporated into the seed pool before proceeding to the final-jailbreak phase. In the final-jailbreak phase, alongside Role-play and Contextualization, we introduce the Expand mutation operator. This operator enhances efficiency by adapting successful jailbreaking templates to address new questions.
@@
We present the system prompts for the three mutators designed for the LLM helper in the highlighted box. Examples of the “Role-play", “Contextualization", and “Expand" mutators are shown in Table 11 (appendix).
Optimization schemes. When optimizing the jailbreaking templates, we consider both semantic coherence and prompt length.
@@ 4.3 Two-level Judge Module
To evaluate whether a response is jailbroken, the judge model should be based on its harmfulness and its relevance to the harmful question. In our method, we implement a two-level judge module to distinguish genuine successful jailbreaks.
@@
The first judge model is a fine-tuned RoBERTa model using manually labeled datasets. Follow [55], the fine-tuning was conducted with a batch size of 16, a learning rate of 1e-5, and a maximum sequence length of 512, using the Adam optimizer with a linear decay schedule and a 10% warm-up ratio. However, this model could only identify responses containing illegal information without considering whether the response addressed the harmful query. This led to inconsistencies between the response and the query.
@@
To address this, we further used GPT-3.5 turbo for an additional round of judgment. This additional check improved the credibility of the jailbreak assessment and ensured the response content aligned with the query. The system prompt for this second judge model is provided in the highlighted box (appendix). This prompt guided GPT-3.5 turbo in assessing the jailbreak status and ensuring the response matched the query. We instructed the LLM to output a score from 1 to 10, with 10 indicating a fully jailbroken response. Upon analyzing responses corresponding to different scores, we observed that once the score reaches 4, the responses contain malicious and misleading content. However, when the score is below 7, even though most of the generated content is malicious, its relevance to the harmful questions is sometimes weak, as shown in Table 12 (appendix). Once the score reaches 8, most responses not only successfully generate jailbreak results but also closely match the harmful questions. Therefore, in our experiments, we set the default jailbreak score threshold to 8 for the ChatGPT-based judge model.

### v4

@@ 29                   $S\leftarrow S\cup\{T\}$
Algorithm 1 Jailbreak Prompts Generation
@@ 4.2 Template Mutation and Optimization
Mutation schemes. Unlike existing methods that demand extensive manual effort and struggle with generalization, Papillon uses a LLM (e.g., GPT-3.5 turbo), as a mutator to automatically create jailbreaking templates for specific questions.
In the pre-jailbreak stage, we design Role-play and Contextualization mutation operators to create customized jailbreaking templates tailored to each question. Successful templates from this phase are incorporated into the seed pool before proceeding to the final-jailbreak phase. In the final-jailbreak phase, alongside Role-play and Contextualization, we introduce the Expand mutation operator. This operator enhances efficiency by adapting successful jailbreaking templates to address new questions. Note that we define high-level mutation strategies, with mutators handling the complex evolution of prompts for given questions. Although the high-level framework relies on human knowledge to guide mutators, this reliance remains minimal.
@@
We present the system prompts for the three mutators designed for the LLM helper in the highlighted box.
Optimization schemes. When optimizing the jailbreaking templates, we control both semantic coherence and prompt length through system prompt instructions and the hyperparameter of the output length respectively.
@@ 4.3 Two-level Judge Module
To evaluate whether a response is jailbroken, the judge model should be based on its harmfulness and its relevance to the harmful question. In Papillon, we implement a two-level judge module to distinguish genuine successful jailbreaks.
@@
The first judge model is a fine-tuned RoBERTa model using manually labeled datasets. Following [54], the fine-tuning was conducted with a batch size of 16, a learning rate of 1e-5, and a maximum sequence length of 512, using the Adam optimizer with a linear decay schedule and a 10% warm-up ratio. This model is used to identify responses containing illegal information without considering whether the response addressed the harmful query.
@@
We further used GPT-3.5 turbo for an additional round of judgment. This additional check improved the credibility of the jailbreak assessment and ensured the response content aligned with the query. While GPT can also be used for illegal content detection, we employ RoBERTa as an initial step to reduce the query costs of LLMs. Only outputs flagged as malicious by RoBERTa are further evaluated by GPT. The system prompt for this second judge model is provided in the highlighted box (appendix). This prompt guided GPT-3.5 turbo in assessing the jailbreak status and ensuring the response matched the query. We instructed the LLM to output a score from 1 to 10, with 10 indicating a fully jailbroken response. Upon analyzing responses corresponding to different scores, we observed that once the score reaches 4, the responses contain malicious and misleading content. However, when the score is below 7, even though most of the generated content is malicious, its relevance to the harmful questions is sometimes weak.Once the score reaches 8, most responses not only successfully generate jailbreak results but also closely match the harmful questions. Therefore, in our experiments, we set the default jailbreak score threshold to 8 for the ChatGPT-based judge model.
5 Experiments
5.1 Experimental Setup
Experimental metrics. We evaluate the effectiveness of Papillon through two metrics, i.e., Attack Success Rate (ASR) and Average Queries (AQ). To make a fair comparison, we calculate the AQ only for successful jailbroken questions. AQ can represent the query overhead of different attacks. Note that both Papillon and baseline attacks use GPT-3.5 turbo in prompt generation and jailbreak judgment procedures, but we use GPT-4 to evaluate attack performance in terms of the output’s relevance to the harmful query for a fairer and more accurate comparison. The length of the generated prompts is controlled using the $max\_new\_tokens$ hyperparameter. Specifically, for the “role-play" and “contextualization" mutations, this value is set to 200 tokens, while an additional 100 tokens are allocated for the “expand" mutation.
More details of the target LLMs and metrics are shown in the appendix.
Our experiments were conducted on a server equipped with 2 NVIDIA A100 PCIe GPUs, each with 40GB of memory. The server’s CPU is an Intel(R) Xeon(R) Silver 4210R with 10 cores, endowed with 188GB of memory. In terms of software, the server runs on the Ubuntu 22.04.1 LTS operating system. The experiments utilized Python version 3.10.13, CUDA version 12.1, PyTorch version 2.1.2, and the transformers library version 4.36.2.
5.2 Comparison with Baselines
We first compare Papillon with five state-of-the-art jailbreaking attacks, including AutoDAN [30], GCG [62], PAIR [7], Gptfuzzer [54], and TAP [34]. The comparison results are shown in Table 10.
We can see that Papillon can achieve the highest attack success rate compared to the baselines across all the victim LLMs. In terms of open-source models, Papillon can achieve ASRs of 100%, 58%, 100%, and 98% for Vicuna-7B-v1.3, Llama-2-7B-chat, Baichuan2-7B-chat, and Guanaco-7B, respectively, while the highest ASRs of the five baselines only achieve 90% (Gptfuzzer), 52% (TAP), and 58% (GCG), and 48% (TAP), respectively. For proprietary models, Papillon achieves an ASR of 90% for GPT-3.5 turbo, 80% for GPT-4, and 82% for Gemini-Pro, significantly outperforming the baselines. Notably, for GPT-4 and Gemini-Pro, Papillon surpasses the baselines by nearly 60%. Additionally, Papillon requires fewer than 25 query times in most cases to achieve the attack goal.
Since Papillon does not rely on any pre-existing jailbreaking templates to initiate the attack, these results demonstrate its effectiveness as a powerful jailbreaking method. The success of Papillon can be attributed to its fuzz testing framework, mutator, and judge module designs.
5.3 Ablation Studies
Impact of different mutators. In Papillon, we design three mutators: Role-play, Contextualization, and Expand. Unlike existing works that rely on manual effort to design them, we propose novel LLM-empowered mutation modules. To explore their effectiveness, we test the ASR and AQ when Papillon adopts different combinations of them, i.e., “Role-play", “Role-play and Contextualization", and all three mutators. The results are shown in Table 11 (appendix).
As the number of mutator types increases, Papillon achieves a higher ASR while still maintaining a low query number. For example, when targeting GPT-3.5 turbo, Papillon attains an ASR of 76% and an AQ of 17.66 when only using the Role-play mutator. However, it achieves an ASR of 86% and an AQ of 16 when the Scenario mutator is added. When using all the mutators, Papillon reaches an ASR of 90% and an AQ of 22.93.
Impact of pre-jailbreak. Since Papillon starts without existing jailbreaking templates, early queries are challenging due to limited seed availability. To address this, we introduce a pre-jailbreak phase with initial jailbreaking attempts (e.g., 5 times) for each problem. Papillon operates in two stages: pre-jailbreak and final-jailbreak. The impact of pre-jailbreak on Papillon is shown in Table 12 (appendix).
Impact of different judges. We adopt a two-level judge module to determine whether a jailbreak prompt is successful. In this section, we explore the impact of each judge on the attack performance of Papillon. The results are shown in Table 13 (appendix).
The results indicate that the stricter the judge module, the more powerful the jailbreak prompts Papillon can generate to deceive the LLM into producing harmful content. Using only RoBERTa, which assesses whether the generated content is harmful but cannot verify its correlation to the question, may lead to ineffective jailbreak prompts when calculating the attack success rate. A jailbreak prompt will only be considered successful if it is both harmful and relevant to the question.
Impact of prompt length. Unlike existing works that overlook prompt length, we limit the prompt length to 200 tokens in Role-play and Contextualization and 100 tokens in Expand in Papillon to decrease attack costs and improve evasiveness. In this section, we explore the impact of prompt length on attack performance. We vary the prompt length from 50 to 300 tokens in Role-play and Contextualization, and the results are shown in Table 15 (appendix).
As the prompt length increases, the attack success rate improves, making the attack more effective. However, this also incurs higher costs since LLM API pricing is based on token length. This creates a trade-off between attack effectiveness and query cost. In Papillon, we set the default prompt length to 200 tokens. Even with a budget of 50 tokens, we achieve an ASR of 100% for Vicuna-7B-v1.3, 98% for Baichuan2-7B, 60% for GPT-3.5 turbo, and 46% for Gemini-Pro. Note that attackers can dynamically adjust the prompt length when using Papillon to launch the attack.
Table 2: Impact of the query budget of the pre-jailbreak.
Target model Metrics budget=5 budget=10 budget=15 budget=20
Vicuna-7B-v1.3 ASR ($\uparrow$) 100% 100% 100% 100%
AQ ($\downarrow$) 4.70 4.40 5.00 5.80
Llama-2-7B-chat ASR ($\uparrow$) 56% 58% 64% 56%
AQ ($\downarrow$) 40.86 37.41 42.75 31.75
Baichuan2-7B ASR ($\uparrow$) 100% 100% 100% 100%
AQ ($\downarrow$) 5.56 3.90 3.78 5.16
Guanaco-7B ASR ($\uparrow$) 96% 98% 96% 98%
AQ ($\downarrow$) 14.62 7.67 10.79 8.62
GPT-3.5 turbo ASR ($\uparrow$) 88% 90% 90% 94%
AQ($\downarrow$) 16.76 18.04 18.27 16.70
GPT-4 ASR ($\uparrow$) 74% 80% 86% 82%
AQ($\downarrow$) 37.08 27.20 23.09 32.90
Gemini-Pro ASR ($\uparrow$) 80% 82% 84% 96%
AQ($\downarrow$) 21.86 19.02 21.18 16.56
Table 3: Mutation with smaller models.
Target model Metrics Vicuna-7b-v1.3 Guanaco-7B
Vicuna-7B-v1.3 ASR ($\uparrow$) 100% 100%
AQ($\downarrow$) 2.86 6.69
Llama-2-7B-chat ASR ($\uparrow$) 58% 56%
AQ($\downarrow$) 36.97 34.68
Baichuan2-7B ASR ($\uparrow$) 100% 100%
AQ($\downarrow$) 3.16 6.30
Guanaco-7B ASR ($\uparrow$) 100% 100%
AQ($\downarrow$) 5.84 12.46
GPT-3.5 turbo ASR ($\uparrow$) 90% 92%
AQ($\downarrow$) 24.43 20.24
GPT-4 ASR ($\uparrow$) 80% 80%
AQ($\downarrow$) 31.28 29.47
Gemini-Pro ASR ($\uparrow$) 76% 74%
AQ($\downarrow$) 24.32 24.19
Table 4: Performance of Papillon when targeting larger open-source LLMs.
Target model ASR AQ
Vicuna-13B-v1.3 100% 12.36
Vicuna-33B-v1.3 100% 4.58
Llama-2-70B 100% 7.20
Llama-3.3-70B-Instruct 64% 30.81
Impact of different query budgets of pre-jailbreak. We investigate the effect of varying query budgets in the pre-jailbreak phase. Specifically, we examine query budgets ranging from 5 to 20 queries, with the results presented in Table 2. It is shown that increasing the query budget in the pre-jailbreak phase leads to a rise in the ASR of Papillon, but the improvement becomes less pronounced with higher budgets. For our experiments, we set the query budget for the pre-jailbreak phase to 10 by default before performing the final-jailbreak. Note that the attacker can choose the most suitable query budget according to the attack performance.
Impact of total query budget. We explore the impact of the query budget in Table 14 (appendix). While the attack success rate generally improves with a higher query budget due to the increased number of attempts to mislead the LLM, it does not continue to increase indefinitely. Some jailbreak prompts can still be circumvented by the LLM, even after more than 150 queries. Consistent with existing studies, we set the default query budget to 100. If prompts are ineffective after 100 queries, they are considered unsuccessful.
Mutation with smaller model. By default, we use GPT-3.5 Turbo for high-quality prompt generation. To evaluate the performance of Papillon with smaller models, we tested Vicuna-7B-v1.3 and Guanaco-7B, as shown in Table 3. While Papillon does not strictly rely on large LLMs, smaller models can lower costs, albeit with potential trade-offs in prompt quality and attack success rates. For instance, Papillon achieves a 100% ASR when targeting the proprietary Gemini-Pro model with GPT-3.5 Turbo, whereas using Guanaco-7B results in a reduced ASR of 74%.
Attack against larger open-sourced LLMs. Apart from the four open-sourced models, i.e., LLaMA-2-7B-chat, Vicuna-7B-v1.3, Baichuan2-7B, and Guanaco-7B, we also explore the effectiveness of Papillon against larger models, including Vicuna-13B-v1.3161616https://huggingface.co/lmsys/vicuna-13b-v1.3, Vicuna-33B-v1.3171717https://huggingface.co/lmsys/vicuna-33b-v1.3, and Llama-2-70B [^18] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf, and Llama-3.3-70B-Instruct [^19] https://build.nvidia.com/meta/llama-3_3-70b-instruct. The results are shown in Table 4. We can see that Papillon maintains a high attack success rate (even 100%) with a limited query budget, even when targeting larger LLMs, highlighting the robustness and power of Papillon.
5.4 Attack Transferability
We further investigate the transferability of Papillon. Transferability refers to the extent to which prompts designed to jailbreak the target model can successfully jailbreak another model.
LLM transferability. We first evaluate transferability by using jailbreak prompts generated by the source model to target another LLM. The results are shown in Table 17 (appendix).
It is evident that prompts generated by one model can sometimes successfully mislead other LLMs into generating malicious content. For example, jailbreak prompts generated by Baichuan2-7B achieve a 70% attack success rate against Vicuna-7B-chat. Overall, Papillon demonstrates notable attack transferability across different models.
MLLM transferability. We then assess the transferability of the PAPILLON-generated jailbreak prompts to multimodal large language models (MLLMs). Specifically, we evaluate two commonly used MLLMs, i.e., Llava-1.5-7B and Llava-1.5-13B. The models are provided with an image and a text prompt as inputs, generating text responses. In the experiments, successful jailbreaking templates are printed on white-background images and used as image inputs, while harmful questions serve as text prompt inputs to mislead the MLLMs into generating malicious or harmful responses. We test 50 jailbreak prompts, and the results are presented in Table 5. Notably, the successful prompts for the LLM are also effective in misleading the MLLMs in many instances, achieving an attack success rate of up to 76% for Llava-1.5-7b and 66% for Llava-1.5-13b.
6 Resiliency against Defenses
6.1 Perplexity Filter
Perplexity is typically low for benign prompts but tends to be high for jailbreak prompts, making it useful for detecting jailbreaking attacks. To reduce false negatives and false positives, Alon et al. [1] proposed a classifier based on two features: perplexity and token sequence length. Jailbreak prompts are typically lengthy, often including long suffixes or prefixes, whereas many regular prompts with high perplexity are notably short. As shown in Table 16 (appendix), when applying the perplexity filter [1] to Papillon, we observe that Papillon maintains a high attack success rate. The ASR drop of Papillon is less than 10% in nearly all cases. The potential reason is that the jailbreak prompts generated by Papillon maintain semantic coherence with low perplexity. This allows these jailbreak prompts to bypass the perplexity-based detection mechanism by closely mimicking regular, benign prompts, thus evading detection.
6.2 SmoothLLM
As shown in Table 16 (appendix), after integrating Smooth-LLM into Papillon, the high attack success rate is still high with only a slight drop. The potential reason is that Papillon is designed to be based on semantics rather than simple character matching. As a result, straightforward character replacement is ineffective at preventing Papillon.
Table 5: Transferability of the Papillon-generated jailbreak prompts to the multi-modal LLMs.
Source model Llava-1.5-7B Llava-1.5-13B
Vicuna-7B-v1.3 72% 66%
Vicuna-13B-v1.3 76% 64%
Llama-2-70B 76% 62%
6.3 Llama Guard
Llama Guard [17] is a safeguard model based on the Llama-2-7B architecture, designed to classify prompts and outputs in AI-human interactions as “safe" or “unsafe". Unsafe prompts are filtered to defend against jailbreak attempts. In our study, we targeted LLMs equipped with Llama Guard, with results presented in Table 6. The experimental findings reveal that we can achieve a high attack success rate even under Llama Guard’s defense. This is attributed to our ability to craft more complex and mutated prompts, which Llama Guard misclassifies as “safe," enabling us to bypass the safeguard.
6.4 Hybrid Defense
In addition to single defenses, we also consider two combinations of multiple defenses, referred to as hybrid defenses. More specifically, we consider two kinds of hybrid defenses, i.e., Perplexity filter + SmoothLLM and SmoothLLM + Llama Guard. In the former case, the defender initially employs a perplexity filter to screen prompts. Prompts that pass this filter are then subjected to the SmoothLLM defense. In the latter case, the defender first leverages Llama Guard to filter unsafe prompts and then applies the SmoothLLM defense to the prompts that pass the Llama Guard filter. If the output generated by the model after applying these hybrid defenses still meets the attack criteria, it is considered a successful jailbreak. The results are shown in Table 7.
Hybrid defenses generally provide stronger protection than single defenses. However, Papillon still achieves a high attack success rate in such cases. This success is due to Papillon’s reliance on its semantic prompts with low perplexity, rather than simplistic character matching. These semantic prompts are not only complex but also undergo deliberate mutations, allowing Papillon to generate highly dynamic and contextually adaptive inputs that effectively bypass detection mechanisms.
Table 6: Apply Llama Guard defense to Papillon.
Target model Papillon After Llama Guard
Vicuna-7B-v1.3 100% 96%
Llama-2-7B-chat 58% 58%
Baichuan2-7B 100% 98%
Guanaco-7B 98% 96%
GPT-3.5 turbo 90% 64%
GPT-4 80% 64%
Table 7: Apply hybird defense to Papillon.
Target model Papillon Perplexity + SmoothLLM SmoothLLM + Llama Guard
Vicuna-7B-v1.3 100% 86% 100%
Llama-2-7B-chat 58% 56% 54%
Baichuan2-7B 100% 92% 98%
Guanaco-7B 98% 86% 84%
GPT-3.5 turbo 90% 80% 58%
GPT-4 80% 80% 52%
7 Human Evaluation
Threshold hyperparameter. To validate the selection of the threshold hyperparameter in the LLM-based judge module, we conducted a human evaluation on a subset of the AdvBench Dataset. We generate jailbreak prompts with different scores under the Papillon framework. A total of 30 volunteers participated in this evaluation by providing binary feedback on whether a given prompt is a successful jailbreaking instance or not. The results are summarized in Table 8. We observed that for a score of 7, a generated prompt was still misaligned with human evaluation. For a score of 8 or higher, all outputs were consistently aligned with human evaluation, achieving a 100% alignment rate. This indicates that a score of 8 serves as an appropriate threshold for distinguishing jailbroken outputs that meet human evaluation standards. Based on these findings, we adopt a threshold value of 8 for our experiments.
Prompts quality. To evaluate the interpretability and naturalness of the generated prompts, we developed two criteria for quantifiable human evaluation. For interpretability, we measure how easily humans can understand the mechanism behind the jailbreak prompt. In simple terms, it assesses whether the evaluator can easily comprehend why the prompt triggers the jailbreak effect. Naturalness is assessed based on the readability of the prompt. Both criteria are scored on a scale of 1 to 5, with higher scores indicating better performance. A total of 30 volunteers participated in this evaluation. For the evaluation, we selected 50 manually crafted jailbreak prompts from Gptfuzzer [54] and 50 jailbreak prompts generated by Papillon, corresponding to the same set of questions. For fairness, we ensured that the formats of the artificial and generated prompts were consistent. Additionally, we randomized the order of the 50 manual prompts and our 50 generated prompts to avoid regularity that could influence the evaluation process. The results are summarized in Table 9. We observed that our generated jailbreak prompts match or surpass manually crafted ones in interpretability and naturalness. Papillon can achieve a high ASR while maintaining outstanding prompt quality.
Table 8: Results of human evaluation to determine the optimal threshold hyperparameter on the AdvBench subset.
Score 6 7 8 9 10
# of generated prompts (Papillon) 0 3 15 12 139
# of judged as jailbroken prompts (Human) 0 2 15 12 139
Alignment rate - 66% 100% 100% 100%
Table 9: Human evaluation to the quality of manual prompts and Papillon generated prompts.
Quality Interpretability Naturalness
# of judged as jailbroken prompts (Human) 3.98 3.66
# of generated prompts (Papillon) 4.64 4.60
All Prompts 4.31 4.03
8 Conclusion and Future Work
This paper introduces Papillon, a fuzz testing-driven framework designed to jailbreak large language models. Papillon can automatically generate jailbreak prompts with only black-box access to the target LLMs. Unlike existing methods that rely on predefined jailbreaking templates, Papillon begins with an empty seed pool, eliminating the need for any related jailbreaks. To enhance attack performance, we design three novel question-dependent mutation strategies that generate semantically meaningful and concise mutations. We also proposed to use a two-level judge module to distinguish successful jailbreaks. Extensive experiments on 7 representative LLMs demonstrate the effectiveness of Papillon compared to 5 state-of-the-art attacks. Papillon also shows resistance to advanced defenses and exhibits high transferability.
There are various potential areas for future exploration. First, while we designed three effective mutators in Papillon, there is still room to develop additional mutators for the mutator pool, such as rephrasing or using ciphertext. These potential mutators may further improve the attack success rate. Second, we will evaluate other state-of-the-art prompt compression methods, such as [36] and [59], to determine if they can generate more effective jailbreak prompts compared to our LLM-driven prompt shortening techniques. Third, effective defense against Papillon is necessary to reduce the potential risks of such attacks. Fourth, Papillon mainly focuses on jailbreaking attacks using English words and targets only English models. An interesting question is whether Papillon can also be effective for other language LLMs. We will explore this in the future. Finally, our findings reveal a critical limitation in aligned LLMs: they are effective in generating jailbreaking templates that do not necessarily exhibit explicit malicious intent. This raises significant safety concerns, indicating the need for further research on alignment strategies to mitigate such behaviors.
Acknowledgement
This research is supported by the National Research Foundation, Singapore under its Strategic Capability Research Centres Funding Initiative and Infocomm Media Development Authority under its Future Communications Research & Development Programme. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore. Yanjiao Chen’s work was partially supported by Ant Group through CCF-Ant Research Fund. Qian Wang’s work was partially supported by the NSFC under Grants U2441240 (“Ye Qisun” Science Foundation), 62441238 and U21B2018. The first two authors contributed equally to this work. Chen Chen is the corresponding author.
Ethics Considerations
In developing the Papillon framework, we have carefully considered the ethical implications of our research. This research inherently involves the potential generation of prompts that could lead to harmful or offensive content. However, we have adopted several measures to ensure that our findings are handled ethically and responsibly.
Stakeholders and Potential Impact. The key stakeholders involved in our research include LLM developers, users of LLM-powered systems, and the wider public who might be indirectly impacted by misuse of these technologies. The release of the Papillon framework is intended to assist researchers and developers in identifying and addressing vulnerabilities in LLMs, thereby improving their overall security and contributing to safer and more robust AI systems. However, we acknowledge the risk that malicious actors could use Papillon to bypass existing defenses, potentially leading to harmful applications. This risk is further amplified by the possibility that Papillon could be adapted to attacking multimodal LLMs, potentially inducing outputs that violate their intended ethical constraints, such as generating images with harmful content (e.g., CSAM, non-consenual sexual material, etc.). To mitigate these risks, we emphasize the need for responsible use of the framework, explicitly discourage any unethical applications, and provide clear guidelines to minimize the likelihood of misuse.
Responsible Disclosure and Dual-Use Concerns. Generating harmful outputs presents significant ethical challenges. To address this, we setup a restricted environment to generate harmful prompts during experiments, only allowing the research team to access. While the core components of Papillon will be open-sourced, specific prompts used for successful jailbreaks will not be released to prevent misuse. This approach balances transparency and research utility with minimizing risks of abuse. Additionally, we will actively monitor community feedback and adapt dissemination practices as needed to prevent potential misuse.
Protection of Research Team Members. The research team has been mindful of the psychological and ethical implications of working with potentially disturbing content. We have ensured that all team members are aware of the risks and have access to support if needed. We have established protocols for reporting any distress experienced by team members and provided access to mental health resources.
By adhering to these principles, we aim to ensure that our research on jailbreaking attacks remains within ethical boundaries and contributes to the responsible advancement of AI technology. We hope that our work not only highlights vulnerabilities but also serves as a catalyst for developing stronger defenses, ultimately contributing to a safer deployment of LLMs.
Compliance with the Open Science Policy
We are committed to the Open Science Policy and will make our research artifacts, such as code and non-sensitive data, publicly available upon publication. This ensures that our work can be replicated and verified by others in the community. Specifically, the following artifacts will be made publicly available upon acceptance: 1) The full implementation of the Papillon framework, including all components necessary for replication, such as the fuzz-testing module, mutation strategies, and the two-level judge module; 2) Detailed scripts and configuration files for running the experiments and reproducing all results presented in the paper. This includes scripts for evaluating attack success rates, processing intermediate results, and running comparative baselines.
To prevent misuse, we will not publicly release any harmful content, such as specific jailbreak prompts. These prompts, as well as any outputs containing potentially harmful responses, will remain restricted to the research team and will only be shared with verified researchers upon request. Access requests will be subjected to a strict review process to ensure ethical use. Our goal is to balance transparency with responsibility, safeguarding against potential harm while enabling meaningful scientific progress.
Appendix A Victim Large Language Models